what is the ph of a solution prepared by mixing 18ml of 0.277m ch3co2h? Assume that the volume of the solution is additive and that Ka=1.8*10^-5


1
Expert's answer
2020-11-09T14:11:05-0500

for weak acid concentration of H+ will be:


[H+] = "\sqrt{(K_a *C)}"


where C is the concentration of weak acid

[H+] = "\sqrt{(1.8\times 10^{-5} \times 0.277)}"

so [H+] = 1.9 "\times" 10-3


where pH = -log([H+])

pH = 2.72